### v1

@@ 3 Experiments
In this section, we present the experiments conducted to evaluate the performance of our proposed method.
@@ 3.1 Experiment setup
Benchmark datasets. We used three benchmark datasets to carry out a series of shadow removal experiments: ISTD Dataset (Wang et al., 2018), the adjusted ISTD (ISTD+) Dataset (Le and Samaras, 2019), and SRD Dataset (Qu et al., 2017). We use the ISTD dataset, which is composed of 1330 training and 540 testing triplets. Each triplet includes shadow images, masks, and shadow-free images. This dataset provides a robust set of images that serve as the basis for our shadow removal algorithm training and testing. We also use the ISTD+ dataset, an enhanced version of the ISTD dataset. It incorporates an image processing algorithm to mitigate the illumination inconsistency between the shadow and shadow-free images found in the original ISTD dataset. The quantity of triplets in the ISTD+ dataset mirrors that in the ISTD dataset. Lastly, we use the SRD dataset, which includes 2680 training and 408 testing pairs of shadow and shadow-free images. It is important to note that this dataset does not provide ground truth shadow masks. To overcome this limitation, we incorporate the predicted masks furnished by DHAN (Cun, Pun, and Shi 2020) for our testing phase only.
@@
Evaluation metrics. We adopt three evaluation metrics to assess the performance of our shadow removal method. These metrics are applied in line with previous works (Qu et al., 2017; Wang et al., 2018; Hu et al., 2019; Guo et al., 2023; Zhu et al., 2022a). 1) Root Mean Square Error (RMSE): We use RMSE as a quantitative evaluation metric for the shadow removal results. The RMSE is computed in the CIE LAB colour space, which models perceptual uniformity across the luminance and chrominance channels. Lower RMSE values indicate better shadow removal, with the metric assessing errors over corresponding pixels in the prediction output and shadow-free target. 2) Peak Signal-to-Noise Ratio (PSNR): We also employ PSNR to evaluate the performance of different approaches in the RGB colour space. Higher PSNR values represent better results. 3) Structural Similarity (SSIM) (Wang et al., 2004): In addition to RMSE and PSNR, we adopt SSIM to further evaluate the shadow removal performance in the RGB color space. Similar to PSNR, higher SSIM values indicate better performance. Our evaluation is conducted on the original image resolutions and with a resolution of $256256x$ 256, following the approach of previous works (Jin et al., 2021; Zhu et al., 2022a; Fu et al., 2021). We compare our technique with various state-of-the-art methods on the ISTD (Wang et al., 2018), ISTD+ (Le and Samaras, 2019), and SRD (Qu et al., 2017) datasets in both quantitative and qualitative ways.
@@
Implementation details. We’ve developed our shadow removal method using PyTorch 1.8, and it’s been tested on a Linux platform with NVIDIA RTX 2080Ti GPUs. Our training process involves random cropping of images into $256256x$ 256 patches. We’ve set the total training epochs to 500 and mini-batch size to 8 across all datasets. The Adam optimizer (Kingma and Ba, 2014) is used with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=1.0\times 10^{-08}$ and an initial learning rate of $lr_{initial}=1e-4$, which linearly decays to $lr_{final}=1e-6$ in the final $200$ epochs.
@@
Figure 5: presents a side-by-side comparison of our method with state-of-the-art shadow removal methods on the ISTD dataset (Wang et al., 2018). The images are arranged from left to right as follows: the input image, methods by DHAN (Cun et al., 2020), Fu et al. (2021), Zhu et al. (2022b), ShadowFormer (Guo et al., 2023), our method, and the ground truth. Zoom in to see the details.
@@
Figure 6: presents a side-by-side comparison of our method with state-of-the-art shadow removal methods on the SRD dataset (Qu et al., 2017). The images are arranged from left to right as follows: the input image, methods by DHAN (Cun et al., 2020), Fu et al. (2021), Zhu et al. (2022b), ShadowFormer (Guo et al., 2023), our method, and the ground truth. Zoom in to see the details.
@@
Figure 7: Visual examples of error maps resulting from our method. The first column displays the original images with shadows. The second column shows the images with the shadows removed. The third column presents the predicted images with the shadows removed. The fourth column illustrates the error maps for the predictions, with white dots representing the error. The color scale on the right side of the error maps ranges from 0 to 255, indicating the magnitude of the error. Zoom in to see the details.
@@ 3.2 Comparison with State-of-the-Art Methods
In this section, we provide a comparable evaluation of our proposed method with various state-of-the-art (SOTA) shadow removal algorithms. These include one traditional method, namely Guo et al. (2012), and a number of deep learning-based methods such as MaskShadow-GAN (Hu et al., 2019), ST-CGAN (Wang et al., 2018), ARGAN (Ding et al., 2019), DSC (Hu et al., 2020), G2R (Liu et al., 2021), Fu et al. (2021), Jin et al. (2021), DHAN (Cun et al., 2020), Zhu et al. (2022b), SG-ShadowNet (Wan et al., 2022), BMNet (Zhu et al., 2022a), and ShadowFormer (Guo et al., 2023). Given the diversity in experimental settings adopted by prior shadow removal methods, we seek a fair comparison by conducting experiments on two primary settings over the ISTD dataset. These settings follow the methodologies of two recent works, Zhu et al. (2022b) and DHAN (Cun et al., 2020). The first setting involves resizing the results into a resolution of $256\times 256$ for evaluation. The second setting maintains the original image resolutions throughout both the training and testing stages. It is important to note that all shadow removal results from the competing methods are either quoted from the original papers, taken from the previous papers or reproduced using their official implementations. This ensures the integrity and fairness of our comparative analysis.
Table 1: The quantitative results of shadow removal using our models and recent method on ISTD (Wang et al., 2018) datasets. We put “-” to denote those models or results that are not available.
@@
Quantitative Measure. The quantitative results on the testing sets over ISTD, ISTD+, and SRD are presented in Tables (Tab. 1, Tab. 2, and Tab. 3), respectively. Our proposed approach achieves superior quantitative performance across all datasets compared to existing methods, as evidenced by lower RMSE and higher PSNR and SSIM in shadow regions, non-shadow regions, and entire images. These consistent improvements demonstrate the strength of our shadow removal on diverse scenes. Interestingly, some existing methods degrade non-shadow regions post processing according to the non-shadow area (NS) results on the ISTD+ dataset in Tab. 3. By contrast, our approach enhances both shadow and non-shadow areas simultaneously. The ability to improve all image regions underscores the precision of our data-driven shadow manipulation techniques. Our method offers several advantages. It can effectively infer consensus point estimates to produce trace-less results by sampling multiple latent variables from the learned distribution. The proposed mask dissociation loss is designed to address the imbalance in supervision between shadow boundary pixels and pixels inside shadows. It does this by leveraging the contextual correlation between shadow and non-shadow regions. This approach allows for a more accurate and balanced supervision, leading to improved shadow detection and removal. In comparison to the most recent state-of-the-art methods, such as those proposed by (Ding et al., 2019) and Fu et al. (2021), which employed very deep and large backbones like U-Net256 and ResNeXt, our method is more efficient, as shown in Tab. 1. Despite using less than 3% of the number of parameters of these competitive approaches, our method delivers outstanding shadow removal performance on test sets.
Qualitative Measure.
@@
Qualitative results on the ISTD (Fig. 5) and SRD (Fig. 6) datasets further showcase the advantages of our approach compared to other methods. The ISTD examples contain relatively simple scenes with high contextual similarity across images. Despite this, prior arts often generate illumination mismatches and artifacts along shadow boundaries in these samples. By contrast, our method produces results absent of tonal inconsistencies and edge effects. This holds even for subtle shadows in similar contexts that cause issues for existing techniques. We also demonstrate strong generalizability on the more complex and diverse SRD shadows. The ability to attenuate difficult shadows while preserving lighting coherence emphasizes the robustness of our data-driven framework. For instance, Fu et al. (2021) and DHAN (Cun et al., 2020) methods erroneously brighten some areas with poor lightness, such as the fourth row of Fig. 5, resulting in ghosting artifacts. Furthermore, almost all competing methods except ShadowFormer (Guo et al., 2023) fail to maintain the illumination coherencies among shadow and non-shadow areas, which severely disrupts the image structure and patterns, thereby generating sharp boundaries as shown in the fourth and fifth rows of Fig. 5. However, with the benefits of the proposed mask dissociation loss, It is evident that our techniques can effectively balance the supervision between shadow boundary pixels and pixels inside shadows. This improves shadow removal, even in complex lighting conditions and intricate shadow patterns. Furthermore, our approach enhances the overall image quality by preserving the integrity of non-shadow regions, thereby providing a more natural and visually pleasing result. On the contrary, the image structures on the SRD dataset are more complex and often feature various colours. In these SRD dataset samples, previous work has sometimes struggled to accurately remove shadows, particularly in areas with intricate structures or varied colours, such as the blue poster in the third row of Fig. 6. The complexity of the image structures and the diversity of colours can pose significant challenges for shadow removal algorithms, leading to less-than-optimal results. Despite the complexity and colour diversity, our method effectively removes shadows while preserving the original colours and structures. This results in high-quality images that maintain the integrity of the original scene, proving the effectiveness of our approach in handling complex and diverse image data.
@@
Table 2: The quantitative results of shadow removal using our models and recent methods on SRD (Qu et al., 2017) datasets. We put “-” to denote those models or results that are not available.
@@
Table 3: Quantitative comparison of our method with the state-of-the-art methods on ISTD+ datasets (Le and Samaras, 2019).
@@ 3.3 Generalization Ability
To validate the generalization ability of our proposed approach, we conducted experiments on the SBU-Timelapse dataset (Le and Samaras, 2022). This dataset is unpaired, meaning it does not include shadow-free images as ground truth. Therefore, we adopted non-reference evaluation indicators: Gradient Structure Similarity (NRSS) and Natural Image Quality Evaluator (NIQE) (Li et al., 2022). The NRSS was used to assess the naturalism of the created shadow-free images, while the NIQE was used to measure the perceptual similarity. For the NRSS metric, a higher value indicates better performance, meaning the denoised image is more similar to the original noisy image. For the NIQE metric, a lower value is indeed better, indicating that the image has a higher perceptual quality. We compared our method with several state-of-the-art methods that do not require masks during inference, including SID (Le and Samaras, 2022), DHAN (Cun et al., 2020), and SG-ShadowNet (Wan et al., 2022). As shown in Table 4, our approach surpasses the other three approaches across all metrics. When compared to SG-ShadowNet (Wan et al., 2022), our method managed to reduce the NIQE from 7.6532 to 3.7532, and achieved an increase in NRSS from 0.9324 to 0.9632. Our method also exhibited better performance when compared to SID (Le and Samaras, 2022) and DHAN (Cun et al., 2020). As depicted in Figure 8, our method achieves reasonable perceptual performance in complicated environments, thereby demonstrating the generalization ability of our method. These results highlight the effectiveness of our proposed method in handling complex and diverse image data.
Table 4: Quantitative comparison of our method with the state-of-the-art methods on the SBU-Timelapse dataset (Le and Samaras, 2022). $\uparrow$” indicates the higher the better and $\downarrow$” indicates the lower the better.
@@
Figure 8: presents a side-by-side comparison of our method with state-of-the-art shadow removal methods on the SBU-Timelapse dataset (Le and Samaras, 2022). The images are arranged from left to right as follows: the input image, methods by SID (Le and Samaras, 2022), DHAN (Cun et al., 2020), SG-ShadowNet (Wan et al., 2022), our method. It is important to note that the SBU-Timelapse dataset is an unpaired dataset, meaning it does not include shadow-free images as ground truth. This aspect adds an additional layer of complexity to the task of shadow removal and serves as a testament to the effectiveness of the compared methods. Zoom in to see the details.

### v2

@@ 5 Experiments
In this section, we present the experiments conducted to evaluate the performance of our proposed method. The experiments are organized into the following parts: 1) Experiment Setup: We describe the datasets, evaluation metrics, and implementation details used in our experiments. 2) Comparison with State-of-the-Art Methods: We compare our method against existing state-of-the-art approaches to demonstrate its effectiveness. 3) Generalization Ability: We evaluate the ability of our method to generalize across different datasets and scenarios. 4) Computational Efficiency: We analyze the computational performance of our method, including inference speed and resource requirements. 5) Real-World Case Study of FieldNet in Field Robotics: We demonstrate the practical applicability of our method in real-world field robotics applications. 6) Ablation Study: We conduct ablation experiments to analyze the contribution of individual components of our method.
@@ 5.1 Experiment setup
Benchmark datasets. We used three benchmark datasets to carry out a series of shadow removal experiments: ISTD Dataset (Wang et al., , 2018), the adjusted ISTD (ISTD+) Dataset (Le and Samaras, , 2019), and SRD Dataset (Qu et al., , 2017). The ISTD dataset is composed of 1330 training and 540 testing triplets. Each triplet includes shadow images, masks, and shadow-free images. This dataset provides a robust set of images that serve as the basis for our shadow removal algorithm training and testing. The ISTD+ dataset is an enhanced version of the ISTD dataset. It incorporates an image processing algorithm to mitigate the illumination inconsistency between the shadow and shadow-free images found in the original ISTD dataset. The quantity of triplets in the ISTD+ dataset mirrors that in the ISTD dataset. Lastly, the SRD dataset includes 2680 training and 408 testing pairs of shadow and shadow-free images. It is important to note that this dataset does not provide ground truth shadow masks. To overcome this limitation, we incorporate the predicted masks furnished by DHAN (Cun et al., , 2020) for our testing phase only.
@@
Evaluation metrics. We adopt three evaluation metrics to assess the performance of our shadow removal method. These metrics are applied in line with previous works (Qu et al., , 2017; Wang et al., , 2018; Hu et al., , 2019; Guo et al., , 2023; Zhu et al., 2022a, ). 1) Root Mean Square Error (RMSE): We use RMSE as a quantitative evaluation metric for the shadow removal results. The RMSE is computed in the CIE LAB colour space, which models perceptual uniformity across the luminance and chrominance channels. Lower RMSE values indicate better shadow removal, with the metric assessing errors over corresponding pixels in the prediction output and shadow-free target. 2) Peak Signal-to-Noise Ratio (PSNR): We also employ PSNR to evaluate the performance of different approaches in the RGB colour space. Higher PSNR values represent better results. 3) Structural Similarity (SSIM) (Wang et al., , 2004): In addition to RMSE and PSNR, we adopt SSIM to further evaluate the shadow removal performance in the RGB colour space. Similar to PSNR, higher SSIM values indicate better performance. Our evaluation is conducted on the original image resolutions and with a resolution of $256\times 256$, following the approach of previous works (Jin et al., , 2021; Zhu et al., 2022a, ; Fu et al., , 2021). We compare our technique with various state-of-the-art methods on the ISTD (Wang et al., , 2018), ISTD+ (Le and Samaras, , 2019), and SRD (Qu et al., , 2017) datasets in both quantitative and qualitative ways.
@@
Implementation details. We’ve developed our shadow removal method using PyTorch 1.8, and it has been tested on a Linux platform with NVIDIA RTX 2080Ti GPUs. Our training process involves random cropping of images into $256\times 256$ patches. We have set the total training epochs to 500 and the mini-batch size to 8 across all datasets. The Adam optimizer (Kingma and Ba, , 2014) is used with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=1.0\times 10^{-08}$ and an initial learning rate of $lr_{initial}=1e-4$, which linearly decays to $lr_{final}=1e-6$ in the final $200$ epochs.
@@
Figure 5: A side-by-side comparison of our method with state-of-the-art shadow removal techniques on sample images from the ISTD dataset (Wang et al., , 2018). The images are arranged from left to right as follows: the input image, methods by DHAN (Cun et al., , 2020), Fu et al., (2021), Zhu et al., 2022b , ShadowFormer (Guo et al., , 2023), our method, and the Ground Truth (GT). Zoom in to see the details.
@@
Figure 6: A side-by-side comparison of our method with state-of-the-art shadow removal techniques on sample images from the SRD dataset (Qu et al., , 2017). The images are arranged from left to right as follows: the input image, methods by DHAN (Cun et al., , 2020), Fu et al., (2021), Zhu et al., 2022b , ShadowFormer (Guo et al., , 2023), our method, and the Ground Truth (GT). Zoom in to see the details.
@@
Figure 7: Visual examples of error maps resulting from our method. The first column displays the original images with shadows. The second column shows the ground truth images with no shadows. The third column presents the predicted shadow-free images after achieved using FieldNet. The fourth column illustrates the error maps for the shadow removal.The colour scale on the right side of the error maps ranges from 0 to 255, indicating the magnitude of the error. Zoom in to see the details.
Figure 8: A side-by-side comparison of our method with state-of-the-art shadow removal techniques on sample images from the SBU-Timelapse dataset (Le and Samaras, , 2022). The images are arranged from left to right as follows: the input image, methods by SID (Le and Samaras, , 2022), DHAN (Cun et al., , 2020), SG-ShadowNet (Wan et al., , 2022), our method. It is important to note that the SBU-Timelapse dataset is an unpaired dataset, meaning it does not include shadow-free images as ground truth. This aspect adds an additional layer of complexity to the task of shadow removal and serves as a testament to the effectiveness of the compared methods. Zoom in to see the details.
@@ 5.2 Comparison with State-of-the-Art Methods
In this section, we provide a comparable evaluation of our proposed method with various state-of-the-art (SOTA) shadow removal algorithms. These include one traditional method, namely Guo et al., (2012), and a number of deep learning-based methods such as MaskShadow-GAN (Hu et al., , 2019), ST-CGAN (Wang et al., , 2018), ARGAN (Ding et al., , 2019), DSC (Hu et al., , 2020), G2R (Liu et al., , 2021), Fu et al., (2021), Jin et al., (2021), DHAN (Cun et al., , 2020), Zhu et al., 2022b , SG-ShadowNet (Wan et al., , 2022), BMNet (Zhu et al., 2022a, ), and ShadowFormer (Guo et al., , 2023). Given the diversity in experimental settings adopted by prior shadow removal methods, we seek a fair comparison by conducting experiments on two primary settings over the ISTD dataset. These settings follow the methodologies of two recent works, Zhu et al., 2022b and DHAN (Cun et al., , 2020). The first setting involves resizing the results into a resolution of $256\times 256$ for evaluation. The second setting maintains the original image resolutions throughout both the training and testing stages. It is important to note that all shadow removal results from the competing methods are either quoted from the original papers, taken from the previous papers or reproduced using their official implementations. This ensures the integrity and fairness of our comparative analysis.
TABLE II: The quantitative results of shadow removal using our models and recent state-of-the-art methods on ISTD (Wang et al., , 2018) datasets. We put “-” to denote unavailable data.
@@
Quantitative Measure. The quantitative results on the testing sets over ISTD, ISTD+, and SRD are presented in Tables (Tab. II, Tab. IV, and Tab. III), respectively, with generalization validated on SBU-Timelapse (Tab. V). FieldNet ‘Plus’ achieves superior performance across all datasets: ISTD (PSNR: 32.76, RMSE: 4.07), ISTD+ (RMSE: 2.6), SRD (SSIM: 0.981), and SBU-Timelapse (NRSS: 0.9632), outperforming ST-CGAN (ISTD PSNR: 27.44), ARGAN (ISTD RMSE: 5.89), and ShadowFormer (ISTD PSNR: 32.21). Compared to these methods, FieldNet’s lightweight design (2.7M parameters vs. 29.2M for ST-CGAN, 125.8M for ARGAN) ensures efficiency without sacrificing quality. Its unpaired training approach and large synthetic dataset enhance robustness across diverse scenes, unlike paired-data-dependent methods.
Qualitative Measure. Qualitative results on ISTD (Fig. 5) and SRD (Fig. 6) showcase FieldNet’s advantages over ST-CGAN, ARGAN, and ShadowFormer. ST-CGAN and ARGAN often produce illumination mismatches and boundary artifacts (e.g., fourth row, Fig. 5), while FieldNet maintains tonal consistency. ShadowFormer excels in fine detail preservation (e.g., fifth row, Fig. 5), but FieldNet avoids sharp boundaries and ghosting seen in ST-CGAN and ARGAN, as evidenced in Fig. 6’s third row with complex structures. This balance of quality and efficiency highlights FieldNet’s suitability for real-time applications.
@@
The ISTD examples contain relatively simple scenes with high contextual similarity across images. Despite this, prior arts often generate illumination mismatches and artifacts along shadow boundaries in these samples. By contrast, our method produces results absent of tonal inconsistencies and edge effects. This holds even for subtle shadows in similar contexts that cause issues for existing techniques. We also demonstrate strong generalizability on the more complex and diverse SRD shadows. The ability to attenuate difficult shadows while preserving lighting coherence emphasizes the robustness of our data-driven framework. For instance, Fu et al., (2021) and DHAN (Cun et al., , 2020) methods erroneously brighten some areas with poor lightness, such as the fourth row of Fig. 5, resulting in ghosting artifacts. Furthermore, almost all competing methods except ShadowFormer (Guo et al., , 2023) fail to maintain the illumination coherencies among shadow and non-shadow areas, which severely disrupts the image structure and patterns, thereby generating sharp boundaries as shown in the fourth and fifth rows of Fig. 5. However, with the benefits of the proposed mask dissociation loss, It is evident that our techniques can effectively balance the supervision between shadow boundary pixels and pixels inside shadows. This improves shadow removal, even in complex lighting conditions and intricate shadow patterns. Furthermore, our approach enhances the overall image quality by preserving the integrity of non-shadow regions, thereby providing a more natural and visually pleasing result. On the contrary, the image structures on the SRD dataset are more complex and often feature various colours. In these SRD dataset samples, previous work has sometimes struggled to accurately remove shadows, particularly in areas with intricate structures or varied colours, such as the blue poster in the third row of Fig. 6. The complexity of the image structures and the diversity of colours can pose significant challenges for shadow removal algorithms, leading to less-than-optimal results. Despite the complexity and colour diversity, our method effectively removes shadows while preserving the original colours and structures. This results in high-quality images that maintain the integrity of the original scene, proving the effectiveness of our approach in handling complex and diverse image data.
Error Visualization: To further analyze the performance of our method, we present error maps in Fig. 7. These error maps provide a visual representation of the difference between the ground truth (shadow-free image) and the output of FieldNet. The first column of Fig. 7 displays the original images with shadows. The second column shows the ground truth images with no shadow. The third column presents the predicted shadow-free images after our shadow-removal process. The most important information lies in the fourth column, which illustrates the error maps for the predictions. Coloured areas in these error maps represent the pixels where the largest discrepancies between the ground truth and the prediction occur. The colour scale on the right side of the error maps ranges from 0 to 255, with higher values indicating a larger error (i.e., a greater difference between the predicted shadow-free image and the ground truth). By examining these error maps, we can gain valuable insights into the strengths and weaknesses of our approach.
@@
TABLE III: The quantitative results of shadow removal using our models and recent state-of-the-art methods on SRD (Qu et al., , 2017) datasets. We put “-” to denote unavailable data.
@@
TABLE IV: The quantitative results of shadow removal using our models and recent state-of-the-art methods on ISTD+ datasets (Le and Samaras, 2019).
@@ 5.3 Generalization Ability
To validate the generalization ability of our proposed approach, we conducted experiments on the SBU-Timelapse dataset (Le and Samaras, , 2022). This dataset is unpaired, meaning it does not include shadow-free images as ground truth. Therefore, we adopted non-reference evaluation metrics: Gradient Structure Similarity (NRSS) and Natural Image Quality Evaluator (NIQE) (Li et al., , 2022).
  • NRSS: Measures the naturalism of the generated shadow-free images. A higher NRSS value indicates better performance, meaning the generated image is more similar to a natural image.
  • NIQE: Evaluates perceptual quality. A lower NIQE value indicates better perceptual quality, meaning the image appears more natural to human observers.
We compared our method with several state-of-the-art methods that do not require masks during inference, including SID (Le and Samaras, , 2022), DHAN (Cun et al., , 2020), and SG-ShadowNet (Wan et al., , 2022). As shown in Table V, our approach surpasses the other methods across all metrics. Specifically:
  • Compared to SG-ShadowNet, our method reduced the NIQE from 7.6532 to 3.7532 and increased the NRSS from 0.9324 to 0.9632.
  • Our method also outperformed SID and DHAN, demonstrating superior generalization ability in handling diverse and complex image data.
Figure 8 shows that our method achieves reasonable perceptual performance in challenging environments. While some methods, such as ShadowFormer in Figures 5 and 6, may produce visually detailed results that appear closer to the ground truth in certain cases, our method emphasizes global illumination consistency, which is particularly effective in complex lighting conditions. This trade-off allows our method to achieve superior quantitative performance while maintaining high perceptual quality, as evidenced by the NRSS and NIQE metrics.
These results highlight the effectiveness of our proposed method in handling complex and diverse image data, demonstrating its strong generalization ability. In future work, we aim to improve the preservation of fine-grained details further while maintaining global illumination consistency.
TABLE V: Quantitative comparison of our method with the state-of-the-art methods on the SBU-Timelapse dataset (Le and Samaras, , 2022). $\uparrow$” indicates the higher the better and $\downarrow$” indicates the lower the better.
@@
To further enhance FieldNet’s generalization across datasets like ISTD, ISTD+, SRD, and SBU-Timelapse, modifications could include integrating real-world paired shadow/shadow-free data to reduce synthetic bias, applying multi-domain augmentations (e.g., urban, indoor scenes) to broaden scene diversity, and implementing adaptive loss weighting to handle extreme lighting conditions. These enhancements, detailed in Sec. 6, aim to address limitations in complex shadow patterns and improve robustness beyond the current synthetic dataset’s scope.